Two trains , each having a speed of 30km/h, are headed at each other on the same straight track .  A bird that can fly 60 km/h flies off the front of one train when they are 60 km apart and heads directly for other train . On reaching the other train , the bird flies directly back to the first train and so forth . What is the total distance the bird travels before the trains collide ?

Grade:12th Pass

5 Answers

aakash
9 years ago
60km infinity
Pushkala Krishnan
9 years ago
pl post the method .
PRAJVAL98
9 years ago
relative velocity of train wrt the other train is 30-(-30)=60km/h
relative distance is 60 km
therefore time taken for collision is 60km/60km/h= 1hr
 
total distance the bird will cover in 1hr = 60km/h * 1h=60km
Anjali
7 years ago
Each train is having a speed of 30km/hr.
Their initial relative distance is 60km.
The bird flies off the front of the first train when both the trains are 60km apart from each other.
Speed of the bird=60km/hr
       Relative distance=Relative speed*Time
        Relative speed is the sum of the speeds of two objects when they are moving towards each other or in opposite directions  and it is the difference of the speeds of the objects when they are moving in the same direction.
Here,
time taken for the bird to meet the second train coming in opposite direction=60km/(30+60)km/hr=60/90hrs=2/3hours
distance travelled by the bird when it meets the second train=2/3hrs*60km/hr=40km
distance travelled by the second train=2/3hrs*30km/hr=20km
So,after the bird reaches the second train ,the bird has travelled 40km and the train 20km.
Distance travelled by the first train in 2/3hrs=2/3hrs*30km/hr=20km
Now the bird sets off from the second train with a speed of 60km/hr and from a distance of 20km from the first train as their relative separation now has become
20km{40km(distance travelled by the bird on reaching the train B)-20km(distance travelled by train A)}
Time taken for the bird to reach back train A=Relative distance(20km)/Relative speed(60km/hr+30km/hr)
                                                                =20/90hrs=2/9hrs
Distance travelled by the bird in the direction of train B=2/9hrs*60km/hr=40/3km
 
Crisp dunk
2 years ago
The answer will be 60km and it'll take the bird infinite number of trips to do that.
 
In its first flight
The bird will meet the train in time t1=2/3hr and travel a distance S1=40km. During this time period the trains would also cover a distance of 20km each, now the remaining distance between both of them is 20km.
 
In second flight
The bird will meet the train in time t2=2/9hr and will travel a distance S2=40/3km. During this time period the trains would also cover a distance of 20/3km each and now the remaining distance between them would be 20/3km.
 
Similarly we will get a patter of infinite trips
t3=2/27hr and S3=40/9km
t4=2/81hr and S4=40/27km
......
This is an infinite GP
Total distance will be
= (40 + 40/3 + 40/9 + 40/27........)km
= 60km